### v1

@@ 5 Multimodal Data Feature Representation
The goal of multimodal data feature representation is to obtain optimal features and data layouts for multimodal data. This process is based on insights from both the data itself and the queries performed on it, aiming to improve the quality of the high-dimensional learned index and ultimately improve query performance. As shown in Fig 5, this process can be divided into two parts: (1)Feature Embedding and Measurement: We obtain representative features of multimodal data using our unique measurement process by concerning intrinsic and extrinsic metrics to select the embedding model with the highest score so that the features can express multimodal data more effectively. (2)Feature Enhancement: We optimize features through hyperspace transformation and movement by analyzing data patterns and learning from query behaviors to adjust feature vectors and overall data layout, making index and rich hybrid queries more efficient.
Figure 5: Overview of Multimodal Data Representation
@@ 5.1 Feature Embedding and Measurement
Figure 6: Workflow of Feature Embedding and Measurement
Feature embedding is the process of converting multimodal raw data into a set of features, while the measurement procedure serves as an assessment to determine the effectiveness of the embedding model, aiding in the selection of the more representative features for subsequent indexing, and querying tasks. We first provide an overall workflow of embedding and measurement, then delve into details. The overall workflow is shown in Fig 6. In the feature embedding process, MQRLD offers an embedding model pool for embedding the given multimodal raw data. The embedded features are then subjected to the measurement process, where we score each model using a data-aware intrinsic metric and a query-aware extrinsic metric to select the embedding model with the highest score.
@@ 5.1.1 Embedding
To initialize, we introduce the construction of the embedding model pool. In the field of feature embedding domain, numerous existing works have proposed diverse techniques for efficiently representing raw data. Table 4 lists some of the notable works that have proven successful in multimodal data feature embedding, including image data, text data, audio data, etc., which we have collected and consolidated into the embedding model pool to support the embedding of wide-ranging attributes of multimodal data. Nevertheless, there is still a critical need to select a suitable model for different scenarios, as the performance of these embedding models can vary significantly depending on factors such as data characteristics and query requirements. For instance, compared to CLIP4Clip, X-CLIP uses temporal embedding models to capture dynamic actions and interactions over time, making it more suitable for the task of embedding features from a given soccer match video. Conversely, CLIP4Clip can effectively capture static, frame-level information, which is more appropriate for life recordings that consist of fragmented information with weak temporal correlations. Consequently, measurement metrics are needed to measure the embedding features, demonstrating the effectiveness of each embedding model.
@@ 5.1.2 Measurement
Evaluating feature embedding models for multimodal data is a challenging task because it requires consideration of generalization ability, fidelity, and support for high-quality downstream query tasks. To address this, we propose a combined measurement metric: widely-used extrinsic metric, complemented by our innovative intrinsic metric, to provide a coarse-to-fine-grained evaluation paradigm. As shown in Fig 6, we define a scoring system to evaluate the effectiveness of each modal, which is formulated as:
@@
where the weights ${w_{1},w_{2},w_{3}}$ are used to control the influence of different metrics on the final measurement. $S_{1}$ is an extrinsic measurement metric, focusing on the performance in downstream tasks, while $S_{2}$ and $S_{3}$ are intrinsic measurement metrics, emphasizing generalization and fidelity ability, respectively.
@@
SC is measured from the perspective of the whole dataset. Features embedded by an effective embedding model from raw data should exhibit well-clustered characteristics, that is, the intra-class distance (distance within each class) should be small and the inter-class distance (distance between different classes) should be large, facilitating feature-related retrieval. To assess the quality of clusters formed by embedding models, we innovatively introduce the SC metric to provide a coarse-grained score that indicates both the consistency within clusters and the separation between clusters. Specifically, we embed the features using $n$ candidate embedding models:
@@
Unlike the global consideration of SC, FID focuses on feature quality for each individual multimodal object. Recent advancements in text-to-image diffusion models have yielded impressive results in generating realistic and diverse images, for example, the remarkable progress in Stable Diffusion makes it possible to generate high-fidelity images with a CLIP text embedding model. Thanks to CLIP, building a bridge between text and image, we could reconstruct the embedded image with a powerful diffusion model, which gives us a chance to have a comprehensive understanding of the visual essence. FID reflects the difference between the original image and the generated image. We calculate the FID of each image’s fine-grained feature distribution and use the FID to judge the fidelity of feature embedding for each multimodal object. Specifically, for embedding model $Model_{j}$, we input the embedded features into the same pre-trained high-performance generative model like Stable Diffusion (Rombach et al., 2022) to get the reconstructed images $\hat{I}_{j}^{{}^{\prime}}$, $\hat{I}_{j}^{{}^{\prime}}=\text{Diffusion}(X_{n})$. Then we input reconstructed images in turn with the original images $\hat{I}_{j}$ into the inception network to get FID value $\text{FID}_{j}=\text{Inception}(\hat{I}_{j},\hat{I}_{j}^{{}^{\prime}})$, and the score of $S_{3}$:
@@
A small FID value between the generated image and the original image indicates that the embedding model has excellent performance and can effectively learn the key information from the original data.
Integrating both coarse-grained and fine-grained measurements with a weighted approach, we derive the final inner score (combination of $S_{2}$ and $S_{3}$). This inner score is then fused with the external score, allowing for a comprehensive selection of the embedding model for different scenarios.
@@
To demonstrate the effectiveness of our measurement method, we select a set of image feature embedding models for evaluation. The experiments utilize a real dataset AI Challenger, with evaluation methods including the SC (Silhouette Coefficient), IN (Intrinsic Measurement), and IN + EX (a combination of Intrinsic Measurement and Extrinsic Measurement). The scoring results are calculated based on different evaluation methods:
@@
In the experiment, we select the image feature embedding models RN50, ViT-B/16, and RN50x64 as experimental benchmarks. Fig 7 illustrates the scoring result of different feature embedding models across various evaluation methods compared with the result of downstream query tasks. The x-axis represents different sizes of image samples and downstream query tasks in the experiment, and the y-axis represents the scoring results after normalization. The scoring results for the downstream query task show that RN50x64 achieves the highest score, followed by ViT-B/16 and RN50, reflecting the actual effectiveness of their respective features in the real query task after embedding. This scoring method effectively distinguishes the performance of different embedding models for downstream query tasks and aligns with the actual scoring results. Comparing the scoring results from different evaluation methods (SC, IN, IN + EX) with those of the downstream query tasks, we find that the SC scoring results show more generalization, as they always fluctuate in a small interval. On the other hand, scoring results IN and IN + EX show more variation and align closely with the scoring results of downstream query tasks. This differentiation arises because SC emphasizes the generalization ability of the embedding model, while IN integrates FID which measures the fidelity of image features, improving evaluation accuracy compared to SC. IN + EX evaluation incorporates feedback from downstream query tasks, providing a more accurate assessment of each embedding model and coming closer to their actual effectiveness.
Figure 7: Comparative Evaluation Results of Different Embedding Models
@@ 5.2 Feature Enhancement
A well-designed representation of multimodal data is crucial for reducing index scan numbers and speeding up queries. This leads to our feature enhancement process that effectively adjusts the feature vectors and data layout by gathering similar multimodal data together, which can improve the quality and speed of the high-dimensional index and ultimately enhance query performance. To be specific, as shown in Fig 8, we treat a high-dimensional feature vector as a data point in hyperspace, and then relocate it to an optimal new location through a unique "projection", considering both the dataset characteristics and query behaviors (Section 5.2.1). The specific method of this projection includes Hyperspace Transformation (Section 5.2.2) and Hyperspace Movement (Section 5.2.3). In the former, the matrix representing the hyperspace transformation is determined, while in the latter, we make further fine-tuning movement. This projection allows the most discriminative dimensions of the data to be emphasized, and similar feature vectors are clustered closely together. Then, we provide evidence that our method successfully improves the performance of clustering within the dataset (Section 5.2.4), which forms a solid foundation for index construction.
Figure 8: An Overview of Feature Enhancement
@@ 5.2.1 Awareness of Projection
There are two aspects we take into account in the projection process, which are the dataset characteristics and query behaviors. From the perspective of dataset, we have the following considerations: (1)Identify and retain information-rich dimensions of certain attributes; (2)Discern underlying distribution of entire dataset; (3)Utilize relationships between multimodal objects.
@@ 5.2.2 Hyperspace Transformation
The hyperspace transformation converts all data points (corresponding to all feature vectors) into new, optimal positions by considering the characteristics of the data and the specific query context. This is achieved through a transformation matrix $T$, which can be derived and optimized in four steps, as detailed below:
@@
All the multimodal objects are managed in DataFrame, from where we select columns of DataFrame to be indexed and represent them with a matrix $D$, as shown in Fig 9, where $D\in\mathbb{R}^{m\times n}$, with each row in matrix $D$ corresponding to a record of a multimodal data object in the DataFrame. $m$ is the total number of records in the DataFrame, and $n$ is the dimension number of all the feature vectors in selected columns.
Figure 9: Matrix D
To enhance the features of multimodal data without further compressing the information and maintain traceability to the original dataset, we define an $N\times N$ transformation matrix $T$ that establishes a one-to-one mapping. Then, we multiply the original matrix $D$ by the $T$ to get enhanced feature vectors denoted as $D_{T}$, where $D_{T}\in\mathbb{R}^{m\times n}$, and $D_{T}=DT$.
@@
By analyzing matrix $D$, we can compute the covariance matrix $C$ ($C\in\mathbb{R}^{n\times n}$) of $D$. The covariance matrix $C$ not only reflects the joint distribution characteristics of the data but also reveals the importance of each dimensions in feature vectors. The $n$ rows of $C$ highlight the distribution pattern of all data records. If the rows of $C$ exhibit small fluctuations, their differences from the mean will be minimal, resulting in a small covariance value. The $n$ columns imply the importance of each dimension and also preserve all dimensional information of the selected columns in the DataFrame, ensuring the specific columns can be accurately mapped using the matrix.
@@
Through the decomposition of $C$, we can construct the transformation matrix $T$ applied to the feature vectors. $T$ comprises a rotation matrix $R$ and a scaling matrix $S$, i.e., $T=RS$. The rotation matrix $R$ indicates the direction of data transformation in hyperspace, while $S$ is a diagonal matrix reflecting the importance of each vector dimension through the scaling of its diagonal elements, thus stretching each dimension of feature vectors. $R$ and $S$ are calculated though eigen decomposition on the covariance matrix $C$ which is represented as:
$C=V\Lambda V^{T}=\begin{bmatrix}v_{1},v_{2},\cdots,v_{n}\end{bmatrix}\begin{bmatrix}\lambda_{1}&0&\cdots&0\\ 0&\lambda_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\lambda_{n}\end{bmatrix}\begin{bmatrix}v_{1}^{T}\\ v_{2}^{T}\\ \vdots\\ v_{n}^{T}\end{bmatrix}$ (7)
where $V=(v_{1},v_{2},\cdots,v_{n})$ is a matrix that is composed of all the eigenvectors of $C$, and $\Lambda=diag(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ is a diagonal matrix, with its diagonal elements being the eigenvalues corresponding to each dimension.
According to the linear transformation relationship of matrices, we know that the eigenvectors in $V$ serve as the basis for the linear transformation of hyperspace coordinates, and the square root of each eigenvalue in $\Lambda$ represents the scaling factor for each dimension in hyperspace. Therefore, we can obtain the rotation matrix $R=V$ and scaling matrix $S=\sqrt{\Lambda}$ from the eigen decomposition results of $C$, and the transformation matrix $T$ can be defined as:
$T=RS=\begin{bmatrix}v_{1},v_{2},\cdots,v_{n}\end{bmatrix}\begin{bmatrix}\sqrt{\lambda_{1}}&0&\cdots&0\\ 0&\sqrt{\lambda_{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sqrt{\lambda_{n}}\end{bmatrix}$ (8)
The invertibility of the hyperspace transformation is an innegligible issue, because after performing indexing, the indexed data needs to be re-transformed into the original feature vector to support query operations on the original data. Therefore, to ensure the reversibility of matrix $T$ and maintain the linearity of data transformation properties, we impose the following constraints on $R$ and $S$:
@@
Constraint (1) indicates that both the rotation matrix $R$ and the scaling matrix $S$ are $n\times n$ square matrices, ensuring that matrix $T$ remains a $n\times n$ square matrix throughout multiply of $R$ and $S$. As a result, after the transformation, the dataset retains $n$ dimensions in the hyperspace with no loss of dimensional information, ensuring that the original dataset is easily regained through one-to-one mapping. Constraint (2) guarantees that $R$ is an orthonormal matrix, with its column vectors forming the standard orthonormal basis of the hyperspace. This ensures that the columns of the matrix transformed by $R$ are independent of other columns, allowing us to focus on specific dimensions corresponding to the feature vectors during query-aware optimization without affecting other dimensions. Constraint (3) specifies that matrix $S$ must be a positive definite matrix, because we only consider scaling the feature vectors in the positive direction. By then, $T$ is calculated and we can get enhanced feature vectors $D_{T}$, where $D_{T}=DT$. These three constraints ensure that matrix $T$ is invertible, which allows us to obtain the original feature vectors directly through the inverse matrix operation by multiplying $D_{T}$ by $T^{-1}$.
@@
We consider the specific requirements of query tasks and adjust the feature vectors to better align with the query context, thereby improving query efficiency and accuracy. To ensure that feature representations accommodate different query contexts, we adjust the hyperspace transformation process by optimizing $T$ through learning from query behaviors. To simplify this optimization process, we decompose the objective of optimizing $T$ into optimizing $R$ and $S$.
To be specific, our goal is to find the optimal $R^{*}$, $S^{*}$ that maximizes the query performance of query workload $Q$, i.e., we formulate this as an optimization problem as
$R^{*},S^{*}=argmax_{Q}QueryScore(R,S)$ (10)
where the $QueryScore$ function returns a scalar score value that evaluates the performance of query execution, with the assumption that a higher score value indicates better performance.
When modeling the $QueryScore$ function, we consider three factors, which are query accuracy, query time, and CBR. This is based on the intuitive idea that an effective index leads to more accurate queries, fewer cross-buckets (resulting in lower I/O), and shorter query times. The detailed description of these three metrics is in Section 4.3.
Based on this consideration, we model $QueryScore$ function:
$QueryScore=\alpha_{1}Score_{acc}-\alpha_{2}Score_{CBR}-\alpha_{3}Score_{time}$ (11)
where $Score_{acc},Score_{CBR},Score_{time}$ are the query scores corresponding to the above metrics of query accuracy, CBR, and query time. The weight value $\alpha=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ is set to standardize each score value and eliminate the impact of different magnitudes among variables.
Given a query workload $Q$, it would be prohibitively costly to directly compute the individual score for each metric. Therefore, we train a random forest model to approximate each score value:
$(Score_{acc},Score_{CBR},Score_{time})=(f_{acc}(R,S),f_{CBR}(R,S),f_{time}(R,S))$ (12)
where $f_{acc}(R,S),f_{CBR}(R,S),f_{time}(R,S)$ are three random forest functions that are used to estimate three score values, which are trained with $R$, $S$, and statistical data from QBS table.
According to equation 10, our goal is to find the optimized matrix $R^{*}$ and $S^{*}$ that maximize the $QueryScore$. The optimization process is presented in Algorithm 1 based on SMAC3 (Lindauer et al., 2022). Matrix $R$ and $S$ are initialized from the original dataset $D$ as described in Step 3, and surrogate model $M$ is obtained using equation 12. Then we iteratively update $S$ using the EI acquisition function and reconstruct $R$ based on updated $S$ to calculate new score value. After $k$ iterations, we return $R^{*}$ and $S^{*}$ that have the highest score value. In this way, we can get optimized feature vectors $D_{T^{*}}=DT^{*}$, where $T^{*}=R^{*}S^{*}$.
0:  dataset matrix $D$, workload $Q$, max iteration k, surrogate model $M$
0:  Optimized matrix $R^{*},S^{*}$
1:  Initialize $R$ and $S$ based on $D$, $Score_{best}=0$, $R_{best}=R$, $S_{best}=S$

### v2

@@ 5 Multimodal Data Feature Representation
Current multimodal data retrieval platforms often fail to fully exploit the synergistic capabilities of the three feature representation stages—embedding, measurement, and enhancement—to optimize data itself for the purpose of enhancing query performance. To fill this gap, our objective in feature representation is to derive the most representative features and optimized data layout for multimodal data, based on insights from both the data itself and the queries executed upon it, to refine the quality of the high-dimensional learned index and thereby ultimately enhance query performance. As shown in Fig 5, this process can be divided into two parts: (1)Feature Embedding and Measurement: We obtain optimal features of multimodal data using our unique measurement process by concerning intrinsic and extrinsic metrics to select the embedding model with the highest score. This serves as the foundation for subsequent optimization of data layout. (2)Feature Enhancement: We optimize overall data layout through hyperspace transformation and movement by analyzing data patterns and learning from query behaviors, making index and rich hybrid queries more effective and efficient.
Figure 5: Overview of multimodal data representation, including feature embedding and measurement, feature enhancement, and optimization based on the query-aware mechanism.
@@ 5.1 Feature Embedding and Measurement
Figure 6: Workflow of feature embedding and measurement. Embedding vector from each model is input into the measurement process consisting of extrinsic and intrinsic metrics to calculate a score, and select the model with the highest score.
Feature embedding is the process of converting multimodal raw data into a set of features, while the measurement procedure serves as an assessment to determine the effectiveness of the embedding model, aiding in the selection of the more representative features for subsequent retrieval tasks. The overall workflow of feature embedding and measurement is shown in Fig 6. In the feature embedding process, MQRLD offers an embedding model pool for embedding the given multimodal raw data. The embedded features are then subjected to the measurement process, where we score each model using a data-aware intrinsic metric and a query-aware extrinsic metric to select the embedding model with the highest score.
@@ 5.1.1 Embedding
In the field of feature embedding domain, numerous existing works have proposed diverse techniques for efficiently representing raw data. Table 4 lists some of the notable works that have proven successful in multimodal data feature embedding, which we have collected and consolidated into the embedding model pool to support the embedding of wide-ranging attributes of multimodal data. Nevertheless, there is still a critical need to select a suitable model for different scenarios, as the performance of these embedding models can vary significantly depending on factors such as data characteristics and query requirements. For instance, compared to CLIP4Clip, X-CLIP uses temporal embedding models to capture dynamic actions and interactions over time, making it more suitable for the task of embedding a soccer match video. Conversely, CLIP4Clip can effectively capture static, frame-level information, which is more appropriate for life recordings that consist of fragmented information with weak temporal correlations. Consequently, measurement metrics are needed to measure the embedding features, demonstrating the effectiveness of each embedding model.
@@ 5.1.2 Measurement
Evaluating feature embedding models for multimodal data is a challenging task because it requires consideration of generalization, fidelity, and support for high-quality downstream query tasks. In this paper, we propose a combined measurement metric: a widely-used extrinsic metric, complemented by our innovative intrinsic metric, to provide a coarse-to-fine-grained evaluation paradigm. As shown in Fig 6, we define a scoring system to evaluate the effectiveness of each modal, which is formulated as:
@@
where the weights ${w_{1},w_{2},w_{3}}$ are used to control the influence of different metrics. $S_{1}$ is an extrinsic measurement metric, focusing on the performance in downstream tasks, while $S_{2}$ and $S_{3}$ are intrinsic measurement metrics, emphasizing generalization and fidelity ability, respectively.
@@
SC is measured from the perspective of the whole dataset. Effective feature embeddings should exhibit well-clustered characteristics, meaning the intra-class distance should be small and the inter-class distance should be large. Therefore, we introduce the SC metric to assess the quality of clusters formed by embedding models. Specifically, we embed the features using $n$ candidate embedding models:
@@
Unlike the global consideration of SC, FID focuses on feature quality for each individual MMO. Recent advancements in text-to-image diffusion models have yielded impressive results in generating realistic and diverse images, building a bridge between text and image. These studies use FID to reflect the difference between the original image and the generated image. Therefore, We use FID to judge the fidelity of feature embedding for each MMO. Specifically, for embedding model $Model_{j}$, we input the embedded features into the same pre-trained high-performance generative model like Stable Diffusion (Rombach et al., 2022) to get the reconstructed images $\hat{I}_{j}^{{}^{\prime}}$, $\hat{I}_{j}^{{}^{\prime}}=\text{Diffusion}(X_{n})$. Then we input reconstructed images in turn with the original images $\hat{I}_{j}$ into the inception network to get FID value $\text{FID}_{j}=\text{Inception}(\hat{I}_{j},\hat{I}_{j}^{{}^{\prime}})$, and the score of $S_{3}$ can be represented as:
@@
A small FID value between the generated image and the original image indicates that the embedding model can effectively learn the key information from the original data.
Integrating both coarse-grained and fine-grained measurements, we derive the final inner score (combination of $S_{2}$ and $S_{3}$). This inner score is then fused with the external score, allowing for a comprehensive selection of the embedding model for different scenarios.
@@
To demonstrate the effectiveness of our measurement method, we select a set of image feature embedding models for evaluation. The experiments utilize a real dataset AI Challenger, with evaluation methods including the SC (Silhouette Coefficient), IN (Intrinsic Measurement), and IN + EX (a combination of Intrinsic Measurement and Extrinsic Measurement). Note that IN itself can function independently in feature measurement, making it applicable for cold start. The scoring results are calculated based on different evaluation methods:
@@
In the experiment, we select the image feature embedding models RN50, ViT-B/16, and RN50x64 as experimental benchmarks. Fig 7 illustrates the scoring result of different feature embedding models across various evaluation methods compared with the result of downstream query tasks. The x-axis represents different sizes of image samples and downstream query tasks in the experiment, and the y-axis represents the scoring results after normalization. The scoring results for the downstream query task show that RN50x64 achieves the highest score, followed by ViT-B/16 and RN50, reflecting the actual effectiveness of their respective features in the real query task after embedding. This scoring method effectively distinguishes the performance of different embedding models for downstream query tasks and aligns with the actual scoring results. Comparing the scoring results from different evaluation methods (SC, IN, IN + EX) with those of the downstream query tasks, we find that the SC scoring results show more generalization, as they always fluctuate in a small interval. On the other hand, scoring results IN and IN + EX show more variation and align closely with the scoring results of downstream query tasks. This differentiation arises because SC emphasizes the generalization ability of the embedding model, while IN integrates FID which measures the fidelity of image features, improving evaluation accuracy compared to SC. Compared to the SC and IN evaluation, we find that IN + EX provides a more accurate assessment of each embedding model, closely reflecting their actual effectiveness. This indicates that the proposed query-ware mechanism, which facilitates the calculation of EX, significantly impacts the optimization of feature representation. Furthermore, although the IN evaluation alone is less precise, it remains capable of selecting relatively suitable embedding models. Therefore, in the absence of a query-aware mechanism (e.g., during platform cold starts), the platform may experience a decline in its capacity but can still maintain normal functionality and deliver reliable performance.
Figure 7: Comparative evaluation results of different embedding models. The combination of IN and EX can best simulate the score value of downstream tasks.
@@ 5.2 Feature Enhancement
A well-designed data layout of multimodal data is crucial for reducing index scans and thus accelerating queries. The primary objective of indexing is to enable effective and efficient data retrieval, and an optimized data layout is fundamental to this process. By physically arranging similar or related data points, such layouts reduce the search scope and significantly enhance indexing performance, particularly in high-dimensional or large-scale scenarios. Currently, most existing approaches, focus on optimizing the data layout through clustering methods. However, these methods typically treat original data as input, neglecting intrinsic data characteristics and their potential impact on layout optimization. In contrast, we propose a novel approach that transforms original data in hyperspace to intrinsically improve the optimization process of obtaining the optimal data layout. To be specific, as shown in Fig 8, we treat a high-dimensional vector as a data point in hyperspace, and then relocate it to an optimal new location through a unique "projection", considering both the dataset characteristics and query behaviors (Section 5.2.1). The specific method of this projection includes Hyperspace Transformation (Section 5.2.2) and Hyperspace Movement (Section 5.2.3). This projection allows the most discriminative dimensions of the data to be emphasized, and similar feature vectors are clustered closely together. Then, we provide evidence (Section 5.2.4) demonstrating that our feature representation process significantly improves clustering performance. This, in turn, validates the effectiveness of our enhancements in optimizing the data layout. Additionally, we show that both learned indexes and vector similarity indexes achieve improved performance when operating on the optimized data layout.
Figure 8: An overview of feature enhancement. Features from Dataframe are represented as a matrix and undergo hyperspace transformation and movement with query-awareness, ultimately being organized into an optimal data layout that exhibits efficient clustering performance.
@@ 5.2.1 Awareness of Projection
There are two aspects we take into account in the projection process, which are the dataset characteristics and query behaviors. From the perspective of dataset, we have the following considerations: (1)Identify and retain information-rich dimensions of certain attributes; (2)Discern underlying distribution of entire dataset; (3)Utilize relationships between MMOs.
@@ 5.2.2 Hyperspace Transformation
The hyperspace transformation converts all data points (corresponding to vectors) into new, optimal positions by considering the data characteristics. This is achieved through a transformation matrix $T$, which can be derived and optimized in four steps, as detailed below:
@@
All the MMOs are managed in DataFrame, from where we select columns of DataFrame to be indexed and represent them with a matrix $D$, as shown in Fig 9, where $D\in\mathbb{R}^{m\times n}$, with each row in matrix $D$ corresponding to an MMO. $m$ is the total number of records in the DataFrame, and $n$ is the dimension number of all selected columns after embedding. To enhance the features of multimodal data without further compressing the information and maintain traceability to the original dataset, we define an $n\times n$ transformation matrix $T$ that establishes a one-to-one mapping. Then, we multiply the original matrix $D$ by the $T$ to get enhanced feature vectors denoted as $D_{T}$, where $D_{T}\in\mathbb{R}^{m\times n}$, and $D_{T}=DT$.
Figure 9: Elements in matrix D correspond to features in DataFrame.
@@
We then compute the covariance matrix $C$ ($C\in\mathbb{R}^{n\times n}$) of $D$. The covariance matrix $C$ not only reflects the joint distribution characteristics of the data but also reveals the importance of each dimensions in feature vectors. The rows of matrix $C$ capture the distribution pattern of data records. Minimal fluctuations in the rows lead to small deviations from the mean, resulting in low covariance. The columns of matrix $C$ represent the importance of each dimension, preserving the complete dimensional information for accurate mapping of the selected DataFrame columns.
@@
Through the decomposition of $C$, we can construct the transformation matrix $T$ applied to the feature vectors. $T$ comprises a rotation matrix $R$ and a scaling matrix $S$, i.e., $T=RS$. The rotation matrix $R$ indicates the direction of data transformation in hyperspace, while $S$ is a diagonal matrix reflecting the importance of each vector dimension through the scaling of its diagonal elements, thus stretching each dimension of feature vectors. $R$ and $S$ are calculated though eigen decomposition on the covariance matrix $C$, $C=V\Lambda V^{T}$.
According to the linear transformation relationship of matrices, we know that the eigenvectors in $V$ serve as the basis for the linear transformation of hyperspace coordinates, and the square root of each eigenvalue in $\Lambda$ represents the scaling factor for each dimension in hyperspace. Therefore, we can obtain the rotation matrix $R=V$ and scaling matrix $S=\sqrt{\Lambda}$ from the eigen decomposition results of $C$, and the transformation matrix $T=RS$.
The invertibility of the hyperspace transformation is an innegligible issue, because after performing indexing, the indexed data needs to be re-transformed into the original vector to support query operations on MMOs. Therefore, to ensure the reversibility of matrix $T$ and maintain the linearity of data transformation properties, we impose the following constraints on $R$ and $S$:
@@
Constraint (1) indicates that after the transformation, the dataset retains $n$ dimensions in the hyperspace with no loss of dimensional information, ensuring that the original dataset is easily regained through one-to-one mapping. Constraint (2) guarantees that $R$ is an orthonormal matrix, ensuring that the columns of the matrix transformed by $R$ are independent of other columns, allowing us to focus on specific dimensions corresponding to the feature vectors during query-aware optimization without affecting other dimensions. Constraint (3) specifies that matrix $S$ must be a positive definite matrix because we only consider scaling the feature vectors in the positive direction. Therefore, after calculating $T$, we can get enhanced feature vectors $D_{T}$, where $D_{T}=DT$. These three constraints ensure that matrix $T$ is invertible through the inverse operation by multiplying $D_{T}$ by $T^{-1}$.
@@
To improve query efficiency and accuracy under different query scenarios, we adjust the hyperspace transformation process by optimizing $T$ through learning from query workloads. To simplify this optimization process, we decompose the objective of optimizing $T$ into optimizing $R$ and $S$.
For a multimodal dataset $D$ and a given query workload $Q$, we aim to achieve high query accuracy, minimal query time, and low CBR when executing queries. Specifically, we seek to identify the optimal $R^{*}$, $S^{*}$ to optimize these objectives as much as possible. Therefore, we formulate this as a multi-objective optimization problem with the objective function:
$\min_{R,S}$ ${\left[f_{\text{time }}(D,q;R,S),f_{CBR}(D,q;R,S),-f_{\text{acc }}(D,q;R,S)\right]}$ (8)
[…]
The constraints of this problem are identical to those in equation (7). In this multi-objective optimization problem, there are three objective functions:
(1) $f_{time}(D,q;R,S)$: Query time, which we aim to minimize.
(2) $f_{CBR}(D,q;R,S)$: CBR, which we aim to minimize.
(3) $f_{acc}(D,q;R,S)$: Query accuracy, which we aim to maximize.
Selecting a multi-objective optimization method involves addressing two key challenges. First, the objective functions are multiple competing black-box objectives, each with high evaluation costs. Second, both $R$ and $S$ are high-dimensional hyperparameters, and their search space is also high-dimensional. To address these challenges, we leverage the Bayesian multi-objective optimization algorithm, MORBO (Daulton et al., 2022). MORBO employs surrogate models to approximate the relationship between parameters and the actual values of the objective functions, thereby reducing evaluation costs. It also designs a coordinated strategy that enables parallel exploration of multiple local regions within the high-dimensional hyperparameter search space to identify global optima. Building on the advantages of the MORBO algorithm, we propose a matrix optimization algorithm, as described in Algorithm 1. First, we initialize multiple trust regions based on the records of Query Accuracy, Query Time, and CBR in the QBS table. Then, within each trust region, we train a local Gaussian process model, select new candidate points, and evaluate new observations. Subsequently, we dynamically update each trust region based on the latest candidate points and observations. Finally, we approximate the Pareto Front (PF) using the new observations and extract the approximate Pareto-optimal solution set P. To obtain unique $R^{*}$ and $S^{*}$ for our feature enhancement process, we assign appropriate weights to the various multi-objective observations based on task-specific requirements. The optimal solution is determined by computing the weighted cumulative sum.
0:  dataset matrix $D$, workload $Q$, objective functions $f=\{f_{acc},f_{time},f_{CBR}\}$, max iteration $k$, trust region parameters:{$n,L_{init},L_{min}$}
0:  approximate Pareto optimal solution set $P$
1:  Initialize the trust regions $T=\{T_{1},...,T_{n}\}$, $X_{0}=\emptyset,Y_{0}=\emptyset$.